OpenStudy (anonymous):

Factor completely: −4x^2 + 16x − 24

OpenStudy (nathan917):

-4(x2-4x+6)

OpenStudy (anonymous):

thnx

OpenStudy (nathan917):

Np.

OpenStudy (anonymous):

-4x^2+16x-24 =-4[x^2-4x+6] =-4[x^2-4x+4-4+6] =-4[(x-2)^2+2] =-4[(x-2)^2-(sqrt2 i)^2] =-4[{x-2+sqrt i}{x-2-sqrt2 i}]